Santiago’s Salsa is in the process of preparing a production cost budget for May. Actual costs in April were:
Santiago’s Salsa
Production Costs
April 2017
24,000
Production Jars of Salsa
Ingredient cost (variable) $19,200
Labor cost (variable) 11,520
Rent (fixed) 5,000
Depreciation (fixed) 6,000
Other (fixed) 1,000
   Total $42,720
Using this information, prepare a budget for May. Assume that production will increase to 28,800 jars of salsa, reflecting an anticipated sales increase related to a new marketing campaign.
Santiago's Salsa
Budgeted Production Costs
May 2017
28,800
Production Jars of Salsa
Ingredient cost $

Labor cost

Rent

Depreciation

Other

Total $

Suppose the wage rate is $12 per hour. How many additional labor hours are needed in May? (Round answer to 0 decimal places, e.g. 5,275.)

Additional labour hours

Calculate the actual cost per unit in April and the budgeted cost per unit in May. (Round answers to 2 decimal places, e.g. 15.20.)

Cost per unit
Actual cost per unit in April $

Budgeted cost per unit in May $
Answer #1
Production calculation cost
Ingredient cost =19200/24000*28800 23040
Labor cost =11520/24000*28800 13824
Rent

5000

5000

Other 1000 1000
Depreciation 6000 6000
Total 48864

Budgeted cost per unit = total cost/no. of units of production

(In may) =48864/28800=$1.7 per unit

Actual cost per unit in April=total cost / no. of units produced

=42720/24000=$1.8 per unit

Actual labor hrs in April =labor cost/cost per hr

=11520/12= 960 hrs

Labor hrs in may =budgeted labor cost/cost per hr

=1152hrs

Additinal required hrs =labor hrs in may - labor hrs in April

=1152-960.

=192 hrs